The St Mary’s Road symbol font

Jeremy Gibbons  Alan Jeffrey

Version 2, March 1994

1 Introduction

This is a brief guide to the St Mary’s Road symbol font, a new symbol font for TEX and L^{A}TEX. It is designed to live with the American Mathematical Society’s fonts, contained in amssymb.sty.

It provides a number of new symbols, including ones for derivation of functional programming (such as ⋎, and ⩕), process algebra ( , □ and ⚡), domain theory (⨅), linear logic (& and ⅋), multisets (⟅x⟆, , and ) and many more. It also fixes some ‘features’ with previous symbols (⊕ used not to be circular, now you can use ⊕ instead) and adds obvious variants of others (such as ↤, ⤇ and ⤆). It is all wrapped up in a L^{A}TEX 2εpackage called stmaryrd, which can be used by saying:

    \usepackage{stmaryrd}

This package understands a large number of options:

• heavycircles says that all of the circular operators such as \oplus and \otimes should by default be heavy, and that \varoplus and \varotimes should refer to the light ones.

• only says that only the symbols listed in the option list should be defined. For example:

         \usepackage[only,mapsfrom,Mapsto,Mapsfrom]{stmaryrd}

says that only the symbols ‘↤’, ‘⤇’ and ‘⤆’ should be defined, which is useful if you use a TEX implementation with limited memory.

2 Symbols

The following operators are defined:

 \Ydown                      \Yleft                      \Yright
 ⅄ \Yup  ⏀ \baro  ⑊ \bbslash
 & \binampersand  ⅋ \bindnasrepma  ⧆ \boxast
 ◫ \boxbar  ⧈ \boxbox  ⧅ \boxbslash
 ⧇ \boxcircle  ⊡ \boxdot  □ \boxempty

⧄ \boxslash                   \curlyveedownarrow          \curlyveeuparrow
\curlywedgedownarrow        \curlywedgeuparrow          \fatbslash
 ⨟ \fatsemi                    \fatslash  ⫴ \interleave
⪦ \leftslice  ⩕ \merge  ⦵ \minuso
\moo                        \nplus  ⌽ \obar
□ \oblong  ⦸ \obslash  ⧁ \ogreaterthan
⧀ \olessthan                  \ovee                       \owedge
⪧ \rightslice  ⫽ \sslash  ⫾ \talloblong
◯\varbigcirc  ⋎ \varcurlyvee  ⋏ \varcurlywedge
⊛ \varoast  ⦶ \varobar  ⍉ \varobslash
⊚ \varocircle  ⊙ \varodot  ⧁ \varogreaterthan
⧀ \varolessthan  ⊖ \varominus  ⊕ \varoplus
⊘ \varoslash  ⊗ \varotimes                  \varovee
\varowedge  ⨉ \vartimes

The following large operators are defined:

□ \bigbox  ⋎ \bigcurlyvee  ⋏ \bigcurlywedge
⫼ \biginterleave              \bignplus  ∥ \bigparallel
⨅ \bigsqcap  ▽ \bigtriangledown  △ \bigtriangleup

The following relations are defined:

\inplus                     \niplus  ⋬ \ntrianglelefteqslant
⋭ \ntrianglerighteqslant ⪿ \subsetplus                 \subsetpluseq
⫀ \supsetplus                 \supsetpluseq  ⊴ \trianglelefteqslant

⊵ \trianglerighteqslant

The following arrows are defined:

⟸\Longmapsfrom  ⟹\Longmapsto  ⤆\Mapsfrom
⤇\Mapsto  ⇽\leftarrowtriangle          \leftrightarroweq
⇿\leftrightarrowtriangle ⚡ \lightning  ⟵\longmapsfrom
↤\mapsfrom  ↗ \nnearrow  ↖ \nnwarrow
⇾\rightarrowtriangle  ⦈ \rrparenthesis  ↓ \shortdownarrow
← \shortleftarrow  → \shortrightarrow  ↑ \shortuparrow
↘ \ssearrow  ↙ \sswarrow

The following delimiters are defined:

⟅ \Lbag  ⟆ \Rbag  ⟅ \lbag
⟦ \llbracket                  \llceil                     \llfloor
⦇ \llparenthesis  ⟆ \rbag  ⟧ \rrbracket
\rrceil                     \rrfloor

Note that \llbracket and \rrbracket are ‘growing’ delimiters that can be used

with \left and \right:

⟦P⟧ ⟦□ P⟧ ⟦ □_{i I}^{a⊕b} P_{i}⟧ ⟦ \begin{matrix}a \\ b \\ c \end{matrix} ⟧ ⟦ \begin{matrix}a \\ b \\ c \\ d \\ e \\ f \end{matrix} ⟧

The following special characters are used in building others:

 \Arrownot                   \Mapsfromchar               \Mapstochar
 \arrownot                   \mapsfromchar

For example, if you type $\Arrownot\Rightarrow$ you get ⇏, and if you type

$\arrownot\rightarrowtriangle$ you get ⇾̸.

Acknowledgements

Thanks to David Murphy for suggestions in the design of the St Mary’s Road font. Thanks to Martin Ward for the first pass of converting the stmaryrd package to L^{A}TEX 2ε. Thanks to Simon Mercer for all the wine at 45 St. Mary’s Road.

Legal rubbish

This document is copyright \begin{matrix}c \\ ◯ \end{matrix} 1991–1994 Alan Jeffrey. The St Mary’s Road fonts are copyright \begin{matrix}c \\ ◯ \end{matrix} 1991–1994 Jeremy Gibbons and Alan Jeffrey. All rights are reserved. The moral right of the authors has been asserted.

You are not allowed to take money for the distribution or use of this file except for a nominal charge for copying, etc.

Redistribution of unchanged files is allowed provided that the whole package is distributed.